# EventHallusion: Diagnosing Event Hallucinations in Video LLMs

### v4

@@ Abstract
Recently, Multimodal Large Language Models (MLLMs) have made significant progress in the video comprehension field. Despite remarkable content reasoning and instruction following capabilities they demonstrated, the hallucination problem of these VideoLLMs is less explored compared with its counterpart in the image domain. To mitigate this gap, we propose EventHallusion, a novel benchmark that focuses on assessing the VideoLLMs’ hallucination toward event, the crux of video analysis. From a hallucination attribution perspective, our EventHallusion benchmark is curated to assess a VideoLLM’s susceptibility toward language priors and vision-language biases. On the other hand, we also propose a simple yet effective method, called Temporal Contrastive Decoding (TCD), to tackle the hallucination problems of VideoLLMs. The proposed TCD method rectifies the model’s bias toward its priors during the decoding stage by comparing the original video with a modified version, in which temporal cues are ambiguited. Through comprehensive evaluation of eight open-source and three closed-source VideoLLMs on the proposed EventHallusion benchmark, we observe that the open-source models suffer significantly from hallucination problems, whereas the closed-source ones perform markedly better. By further equipping open-source VideoLLMs with the proposed TCD approach, evident performance improvements are achieved across most metrics in the EventHallusion benchmark.
@@ 1 Introduction
The field of deep learning has witnessed the extraordinary advancement of Large Language Models (LLMs) [6, 39, 10] and Multimodal Large Language Models (MLLMs) [4, 31, 5, 17, 15]. Benefiting from dramatic semantic reasoning and instruction-following capabilities, LLMs and MLLMs have become versatile assistants across multiple disciplines  [15, 44]. Despite their great success, they are still trapped in the hallucination problem, where models generate plausible but factually incorrect responses [21], potentially threatening reliability and security in applying them in real-world scenarios [36, 16]. To systematically evaluate hallucinations of ImageLLMs, a thread of benchmarks has been developed [25, 12, 32]. Specifically, POPE [25] investigates the hallucination by mining the frequently occurring objects and their combinations. MME [12] and MMBench [32] evaluate hallucinations of the existence, attributes, and relations of objects.
@@
In fact, based on the foundation of ImageLLMs, VideoLLMs have shown significant proficiency in understanding the content of single-frame images [40]. However, their effectiveness diminishes when tasked with comprehending the event, namely the actions performed by humans or objects spanning the entire video [11, 46, 19]. This limitation is further proved by our results presented in Fig.6. To further evaluate the event-oriented hallucinations of VideoLLMs, two recent benchmarks [40, 13] develop distinct event comprehension tasks. Specifically, HallusionBench [13] reverses all frames within a video, subsequently querying the model to determine whether the original event occurs in the altered video. VideoHallucer [40] selects long videos containing multiple events and then queries the model to evaluate their chronological orders. Although advancing the evaluation of event hallucination, they still suffer from the following limitations. On the one hand, the frame-reversed video in HallusionBench can deviate from the distribution of natural videos, as not all events remain meaningful when played in reverse. On the other hand, merely judging the chronological order of events given in the questions fails to probe the model’s authentic ability to comprehend the complex events [^1] A more detailed analysis of event-related VQA samples in HallusionBench [13] and VideoHallucer [40] is provided in the supplementary materials due to page limitations..
When delving into VideoLLMs’ event-related hallucinations, we identify two primary causes rooted in the development processes of prevalent VideoLLMs. Firstly, building upon the foundation of LLMs, VideoLLMs inherently inherit their language priors [3, 2], which manifests as an overconfidence in their language knowledge. As shown in the left column of Fig.1, both Video-LLaVA[28] and ShareGPT4Video[8] are easily misled by the frequently occurring but misleading event “person drink coffee” asked in the question. Moreover, due to the high cost associated with scaling up high-quality video-language data, prevalent VideoLLMs typically acquire multimodal knowledge from large-scale image-language datasets. However, in these datasets, specious event-related captions often arise based solely on a single image, which inevitably introduces spurious scene-event correlations. As shown in the right column of Fig.1, both Video-LLaVA[28] and ShareGPT4Video[8] intuitively make the conclusion of “cats are playing” when perceiving two cats within the video, without respecting the truth.
@@
Additionally, we propose a simple yet effective method called Temporal Contrastive Decoding (TCD) to reduce the event-related hallucinations of existing VideoLLMs in a training-free manner. The core motivation of our TCD approach is to first elicit the spurious priors and then eliminate their adverse effects before making the final decisions. Technically, for each given video, we first construct its counterpart by ambiguiting the temporal cues while preserving the spatial information. This counterpart disentangles the event-unrelated elements that could potentially trigger spurious prior assumptions from the original video. Therefore, we feed both the original video and its counterpart into a VideoLLM and adaptively adjust the prediction distribution of the former using the latter, leading to more reliable final judgments. As a plug-and-play approach, our TCD can consistently boost existing VideoLLM’s performances on the proposed EvenHallusion as shown in Tab.2.
@@
2 Related Works
Video Large Language Models. In the natural language processing field, Large Language Models (LLMs) have demonstrated astounding performance in solving traditional tasks. What’s more, LLMs also exhibit extraordinary generalization in handling self-defined tasks described by user’s instructions, making them capable of playing the role of versatile assistants. Benefiting from such abilities, Multimodal Language Models (MLLMs) have emerged by introducing visual inputs into LLMs, promoting the unified comprehension of both textual and visual content.
Video-supported MLLMs are further proposed since the established vision-text understanding ability of image MLLMs. To enable MLLMs to comprehend video content and the concept of temporality, numerous methods have been put forward. Video-ChatGPT [35] utilizes spatial and temporal pooling operations to conduct video modeling, while Video-LLaMA [47] introduces a Video Q-Former to aggregate frame-level representations. Vista-LLaMA [34] preserves the equal distance between all visual tokens and any language tokens and proposes the temporal Q-Former projector to further enhance the representations of the video and the modeling of temporality. LLaMA-VID [26] proposes a dual-token paradigm, using a context token and a content token to represent each frame, which effectively supports the comprehension of long videos and maintains temporal information. However, despite their excellent performance on existing benchmarks, these methods still suffer the hallucination problem in temporality.
Video Evaluation Benchmarks. Despite the extraordinary advancement of MLLMs, it is of great significance to conduct quantitative evaluations of existing models. To thoroughly evaluate the performance of image MLLMs, a wide scope of efforts has been made [12, 32, 27]. Their evaluation comprises various aspects, including perception, reasoning, and cognition. With the development of VideoLLMs, several benchmarks have also been proposed to assess video comprehension capacities. Tempcompass [33] focuses on the thorough evaluation of the temporal perception ability of VideoLLMs, including action, speed, direction, attribute change, and event order. FunQA [41] proposes utilizing humorous, creative, and magic videos to assess the in-depth comprehension ability of existing VideoLLMs.
Despite previous works sharing similarities with ours, there exist several critical differences. First, we focus on exploring and evaluating the hallucination phenomenon in the event aspects of existing VideoLLMs. Furthermore, we emphasize and validate that the event hallucination phenomenon of video LLMs is influenced by language prior. Second, based on this conclusion, we collect videos that are inclined to cause event hallucinations in VideoLLMs and construct the benchmark for evaluation. Third, we propose a Temporal Contrastive Decoding (TCD) to alleviate the event hallucination, which exhibits effectiveness in multiple VideoLLMs.
@@ 2 Related Works
Hallucination in MLLMs. Hallucination is a phenomenon widely observed in MLLMs. The occurrence of hallucinations is attributed to various factors, including imbalanced training data distribution, weak attention to visual content, and other possible reasons. To alleviate the impact of hallucinations in image MLLMs, several methods have been proposed. Woodpecker [43] adopts a post-hoc correction mechanism that utilizes additional extracted visual cues to rectify previously generated responses, while Volcano [20] proposes a critique-revise-decide process to self-revise their initial generated answers. Besides, LRV-Instruction [30] and HalluciDoctor [45] propose to mitigate this problem from the data perspective to provide more balanced, high-quality instruction tuning datasets.
Some other works propose to alleviate hallucination during the inference phase, in which the contrastive decoding (CD) method has been widely adopted due to its excellent performance and the absence of additional training. VCD [21] employs the original and distorted images for comparison in contrastive decoding, thereby reducing the impact of statistical bias and language priors, while IBD [48] constructs the comparison pair by introducing the image-biased and original MLLMs. Despite the extensive exploration of adopting CD in ImageLLMs, there is still a lack of effort in employing CD to handle hallucinations in VideoLLMs. Inspired by this, we propose Temporal Contrastive Decoding (TCD) to mitigate event hallucinations in VideoLLMs.
@@ 3.1 Data Curation
Susceptibility to vision-language correlation biases. In this category, the primary challenge lies in sourcing the appropriate video data. Since existing publicly available video datasets (e.g, ActivityNet [7], [9], [37]) collect videos from daily life, the events they include cater to the vision-language correlation biases stored in VideoLLMs, resulting in high recognition precision. Therefore, instead of refactoring existing video datasets, we meticulously collect videos from prevalent Internet video platforms and filter them with rigorous cross-validation. Specifically, we begin by manually collecting videos that depict infrequent events based on human consensus. To minimize personal bias, we engage three annotators to assess whether the events in these selected videos are indeed uncommon. Only videos that receive unanimous positive judgments from all annotators are retained as testing samples. Similar to the first category, we also employ the blurring operation on videos to prevent information leakage. We further divide these videos into two groups: (1) Videos where infrequent events occur throughout the entire duration, referred to as “Entire Rare Events” in Fig.2. (2) Videos containing a mix of both frequent and infrequent events, referred to as “Mix Common-Rare Events” in Fig.2.
@@
Binary: “Did the [subject] [event] in the video?”,
@@
“Did the [subject] [event] throughout the entire video?”,
@@
3.2 Evaluation Protocol
Binary question answering. Following previous benchmarks [40, 25], we compare the first word of the model’s answer with the ground-truth answer to determine whether the model answered the question correctly. This approach is justified by the VideoLLM’s exceptional instruction-following capability, which typically enables it to answer binary choice questions with “yes” or “no” at the beginning. We calculate the accuracy as the evaluation metric for binary questions.
Open-ended question answering. For automatic evaluation, we feed both generated and ground-truth answers into GPT-4o[14], using it as a knowledgeable evaluator to determine whether the model answers the question correctly following prior benchmarks [33]. We also use accuracy as the evaluation metric. Detailed templates of prompting GPT-4o are provided in the supplementary materials due to page limitation.
Table 1: Performance comparison of VideoLLMs on EnventHallusion. We exhibit the accuracy of both open-sourced and closed-sourced models on binary classification task and detailed description matching task on our three types of videos respectively. We use red to highlight the best results in close-source VideoLLMs and blue for open-source VideoLLMs. In addition, we show the overall success rate of each model on our binary classification task that answers ”yes” or ”no” in the first word.
@@ 3.2 Evaluation Protocol
VideoChatGPT [35] 14.91 5.50 56.99 3.63 21.57 36.43 4.30
VideoChat2 [24] 16.67 4.59 47.67 1.55 22.55 32.76 2.64
PLLaVA [42] 45.61 16.51 58.55 3.11 81.37 60.64 6.05
LLaMA-VID [26] 30.70 16.51 73.58 7.77 43.14 54.03 10.90
LLaVA-NeXT-Video [22] 49.12 9.17 70.47 4.66 71.57 64.79 6.29
VILA [29] 53.51 20.18 58.55 17.62 83.33 63.33 18.54
ShareGPT4Video [8] 11.40 0.00 67.88 5.18 6.86 49.14 9.82
Video-LLaVA [28] 30.70 8.26 57.51 7.25 41.18 45.97 7.62
Close-source Models
GPT-4V [1] 57.89 11.01 79.27 16.58 92.16 76.53 14.57
Gemini-1.5-Pro [38] 61.40 49.54 83.42 39.90 96.08 80.44 43.38
GPT-4o [1] 83.33 58.72 92.75 41.97 100.00 91.93 48.01
4 Temporal Contrastive Decoding
To mitigate the aforementioned hallucinations in VideoLLMs, the most straightforward solution is to enhance finetuning with additional debiased data. However, this approach is highly resource-intensive, as it requires extensive efforts to collect such videos and annotate corresponding QA pairs. Therefore, we take inspiration from [21] and propose our Temporal Contrastive Decoding (TCD) method, which alleviates event-oriented hallucinations in VideoLLMs by comparing the logits of the original video and its counterpart during the decoding stage.
@@ 4 Temporal Contrastive Decoding
where $S(\cdot)$ represents the temporal downsampling operation. With the generated $z_{t}^{ori}$ and $z_{t}^{con}$, the modulated probability distribution $p_{t}$ at the current timestep can be calculated as:
@@
where $\alpha$ is the hyperparamter to control the contrastive strength during the decoding. Since $z_{t}^{ori}$ and $z_{t}^{con}$ are not always exhibited contrastively, we eliminate the adverse effects of suppressing true positives by directly filtering implausible answers with a dynamic confidence threshold $t$:
@@
$p_{t}[k]=\begin{cases}\operatorname{softmax}\Big{(}z_{t}[k]\Big{)}&\text{if }z_{t}[k]>=t,\\ -\infty&\text{if }z_{t}[k]<0.\end{cases}$ (5)
@@
Table 2: Results of TCD on EventHallusion. Default denotes the standard decoding strategy, while TCD represents our Temporal Contrastive Decoding method. We exhibit the accuracy of the binary classification task. The best results of each category are bolded. “LNV” is short for “LLaVA-Next-Video”.
Model Decoding Entire Mix Misleading Overall
LNV [22] Default 49.12 70.47 71.57 64.79
@@
5 Experiments
In this section, we first evaluate multiple prevalent VideoLLMs on our EventHallusion, including both open-sourced and private MLLMs. Subsequently, we examine the effectiveness of our proposed TCD, and also discuss its more specific design choices.
5.1 Evaluation on EventHallusion
5.1.1 Performances of Human Evaluation
Firstly, to evaluate the quality of our proposed EventHallusion benchmark, we ask three humans to answer the questions and average their resulting scores as the human evaluation results. Meanwhile, to validate the value of evaluating event-oriented hallucinations for VideoLLMs, we also extracted object contents from videos in our EventHallusion benchmark and tested them using both human evaluations and two prevalent VideoLLMs. The above results are shown in Fig.6, where we have the following conclusions: (1) Although prevalent VideoLLMs perform poorly in EventHallusion (i.e, “Event” part in Fig.6), humans can correctly answer almost all questions. This demonstrates that the questions in our EventHallusion benchmark are clear enough to allow for accurate judgments. (2) When recognizing objects in videos, prevalent VideoLLMs can achieve human-level accuracies. However, they significantly lag behind humans when comprehending event-related content. This demonstrates that event-oriented hallucinations confuse VideoLLMs much more severely than object-related ones.
5.1.2 Peformances of Prevalent VideoLLMs
Evaluation Setup. We test 8 open-sourced Video and 3 close-sourced VideoLLMs on our EventHallusion. Open-sourced VideoLLMs comprise VideoChatGPT [35], VideoChat2 [24], PLLaVA [42], LLaMA-VID [26], VideoLaVIT [18], LLaVA-NeXT-Video [22], VILA [29], ShareGPT4Video [8], and Video-LLaVA [28]. For close-sourced VideoLLMs, we select the three most popular video-supported MLLMs, namely GPT-4V [1], GPT-4o[14] and Gemini-1.5-Pro [38], to conduct the systematic comparison of the robustness between open-sourced and close-sourced MLLMs.
For the inference setting of each VideoLLM, we follow their original setting including conversation mode and other hyperparameters. For the number of frames, we sample at least 12 frames for each VideoLLM to ensure the complete preservation of event information. In addition, to ensure the stability and reproduction of our results, we adopt the greedy search as the default decoding strategy for all open-sourced models. For the close-sourced models, we sample 8 frames.
@@ 5.1.2 Peformances of Prevalent VideoLLMs
Performance Analysis. We exhibit the performance of each VideoLLM in Tab 1 on the binary classification and detailed description matching task. As the results indicate, open-source VideoLLMs exhibit a consistent weakness in event hallucinations. In contrast, closed-source VideoLLMs demonstrate greater robustness against event hallucinations across all categories. A deeper analysis is provided below: (1) In evaluating the susceptibility of VideoLLMs to vision-language correlation biases (i.e, “Entire” and “Mix” in Tab.1), most VideoLLMs perform worse when confronted with videos containing a mix of common and rare events, compared to those containing only rare events. This result indicates that frequent events can distract the model’s attention, imposing higher challenges to the model. (2) In evaluating the susceptibility of VideoLLMs to language priors (i.e., “Misleading” in Tab.1), closed-source models demonstrate overwhelming advantages compared to open-source models, suggesting that these closed-source VideoLLMs are built upon language foundation models that are significantly more powerful than their open-source counterparts.
5.2 TCD Evaluation
5.2.1 VideoLLM Baselines.
To evaluate the effectiveness of our TCD in addressing event hallucinations, we adopt LLaVA-NeXT-Video [22], VideoChat2 [24], and VILA [29] as the foundation backbones. LLaVA-NeXT-Video extends the ability of LLaVA [31] based on the projection layer design, while VideoChat2 adopts the Q-Former [23] to bridge the gap between videos and LLM. VILA, on the other hand, employs enhanced visual-language pre-training for better video comprehension. Based on these MLLMs, we explore the performance of TCD across different VideoLLM architectures and designs on EventHallusion.

### v5

@@ Abstract
Recently, Multimodal Large Language Models (MLLMs) have made significant progress in the video comprehension field. Despite remarkable content reasoning and instruction following capabilities they demonstrated, the hallucination problem of these VideoLLMs is less explored compared with its counterpart in the image domain. To mitigate this gap, we propose EventHallusion, a novel benchmark that focuses on assessing the VideoLLMs’ hallucination toward event, the crux of video analysis. From a hallucination attribution perspective, our EventHallusion benchmark is curated to assess a VideoLLM’s susceptibility toward language priors and vision-language biases. On the other hand, we also propose a simple yet effective method, called Temporal Contrastive Decoding (TCD), to tackle the hallucination problems of VideoLLMs. The proposed TCD method rectifies the model’s bias toward its priors during the decoding stage by comparing the original video with a modified version, in which temporal cues are made ambiguous. Through the comprehensive evaluation of eight open-source and three closed-source VideoLLMs on the proposed EventHallusion benchmark, we observe that the open-source models suffer significantly from hallucination problems, whereas the closed-source ones perform markedly better. By further equipping open-source VideoLLMs with the proposed TCD approach, evident performance improvements are achieved across most metrics in the EventHallusion benchmark. Our codes and benchmark data are available at https://github.com/Stevetich/EventHallusion.
@@ Introduction
The field of deep learning has witnessed the extraordinary advancement of Large Language Models (LLMs) (Brown et al. 2020; Touvron et al. 2023; Dettmers et al. 2024) and Multimodal Large Language Models (MLLMs) (Alayrac et al. 2022; Liu et al. 2024a; Bai et al. 2023; Jiao et al. 2024b, a). Benefiting from dramatic semantic reasoning and instruction-following capabilities, LLMs and MLLMs have become versatile assistants across multiple disciplines  (Jiao et al. 2024a; Yin et al. 2023b). Despite their great success, they are still trapped in the hallucination problem, where models generate plausible but factually incorrect responses (Leng et al. 2024), potentially threatening reliability and security in applying them in real-world scenarios (Qian et al. 2024; Jiao et al. 2022). To systematically evaluate hallucinations of ImageLLMs, a thread of benchmarks has been developed (Li et al. 2023b; Fu et al. 2023; Liu et al. 2023b). Specifically, POPE (Li et al. 2023b) investigates the hallucination by mining the frequently occurring objects and their combinations. MME (Fu et al. 2023) and MMBench (Liu et al. 2023b) evaluate hallucinations of the existence, attributes, and relations of objects.
@@
In fact, based on the foundation of ImageLLMs, VideoLLMs have shown significant proficiency in understanding the content of single-frame images (Wang et al. 2024). However, their effectiveness diminishes when tasked with comprehending the event, namely the actions performed by humans or objects spanning the entire video (Du et al. 2024; Zelnik-Manor and Irani 2001; Krishna et al. 2017). This limitation is further proved by our results presented in Fig.6. To further evaluate the event-oriented hallucinations of VideoLLMs, two recent benchmarks (Wang et al. 2024; Guan et al. 2024) develop distinct event comprehension tasks. Specifically, HallusionBench (Guan et al. 2024) reverses all frames within a video, subsequently querying the model to determine whether the original event occurs in the altered video. VideoHallucer (Wang et al. 2024) selects long videos containing multiple events and then queries the model to evaluate their chronological orders. Although advancing the evaluation of event hallucination, they still suffer from the following limitations. On the one hand, the frame-reversed video in HallusionBench can deviate from the distribution of natural videos, as not all events remain meaningful when played in reverse. On the other hand, merely judging the chronological order of events given in the questions fails to probe the model’s authentic ability to comprehend the complex events [^1] A more detailed analysis of event-related VQA samples in HallusionBench (Guan et al. 2024) and VideoHallucer (Wang et al. 2024) is provided in the supplementary materials due to page limitations..
When delving into event-related hallucinations in VideoLLMs, we observe that such errors originate from a consistent issue across different training stages: imbalanced training data distributions, which inject prior knowledge and bias into the model. Initially, building upon LLM backbones, VideoLLMs memorize events with higher frequencies purely through textual statistics, thereby forming strong language-driven priors (Agrawal, Batra, and Parikh 2016; Agarwal, Shetty, and Fritz 2020). For instance, the commonly described event “person drinks coffee” frequently appears in textual corpora, establishing a misleading prior that persists regardless of visual evidence (Fig.1, left). Moreover, due to the high cost associated with scaling up high-quality video-language data, prevalent VideoLLMs typically acquire multimodal knowledge from large-scale image-language datasets. However, similar imbalances surface during the vision-language alignment stage, where frequently co-occurring visual-textual pairs reinforce spurious object-event associations. As shown in the right column of Fig.1, both Video-LLaVA (Lin et al. 2023) and ShareGPT4Video (Chen et al. 2024) intuitively conclude “people are riding the bicycle” simply from the presence of a person and a bicycle in the scene, disregarding the actual video context.
@@
Additionally, we propose a simple yet effective method called Temporal Contrastive Decoding (TCD) to reduce the event-related hallucinations of existing VideoLLMs in a training-free manner. The core motivation of our TCD approach is to first elicit the spurious priors and then eliminate their adverse effects before making the final decisions. Technically, for each given video, we first construct its counterpart by ambiguating the temporal cues while preserving the spatial information. This counterpart disentangles the event-unrelated elements that could potentially trigger spurious prior assumptions from the original video. Therefore, we feed both the original video and its counterpart into a VideoLLM and adaptively adjust the prediction distribution of the former using the latter, leading to more reliable final judgments. As a plug-and-play approach, our TCD can consistently boost existing VideoLLM’s performances on the proposed EventHallusion as shown in Tab.2.
@@
Related Works
Video Large Language Models. Large Language Models (LLMs) have achieved remarkable success in traditional NLP tasks and exhibit strong generalization to user-defined instructions, enabling them to act as versatile assistants. Building on this foundation, Multimodal Language Models (MLLMs) extend LLMs with visual inputs, facilitating unified text-vision understanding. To further incorporate video comprehension and temporality, video-supported MLLMs have emerged. Methods such as Video-ChatGPT (Maaz et al. 2023) apply spatial-temporal pooling, while Video-LLaMA (Zhang, Li, and Bing 2023) introduces a Video Q-Former for frame aggregation. Vista-LLaMA (Ma et al. 2023) enhances temporal modeling via a temporal Q-Former projector, and LLaMA-VID (Li, Wang, and Jia 2023) proposes a dual-token paradigm to represent both context and content per frame. Although these approaches perform well on benchmarks, they still struggle with hallucinations in temporal understanding.
Video Evaluation Benchmarks. Despite the rapid progress of MLLMs, a comprehensive evaluation of their capabilities remains essential. For image MLLMs, various benchmarks have been developed to assess perception, reasoning, and cognition (Fu et al. 2023; Liu et al. 2023b; Li et al. 2024c). With the emergence of VideoLLMs, benchmarks like TempCompass (Liu et al. 2024b) and FunQA (Xie et al. 2023) have been proposed to evaluate temporal perception and complex video comprehension.
Despite previous works sharing similarities with ours, several critical differences persist. First, our work specifically investigates event-related hallucinations in VideoLLMs, attributing their occurrence to language priors and vision-language bias inherited during pretraining. Building on this insight, we curate a benchmark with event-prone videos to facilitate targeted evaluation. Furthermore, we introduce Temporal Contrastive Decoding (TCD), a simple yet effective method to alleviate event hallucinations, demonstrating general improvements across multiple VideoLLMs.
@@
Hallucination in MLLMs. Hallucination is a phenomenon widely observed in MLLMs, arising from factors such as imbalanced training data, weak attention to visual content, and other possible reasons. To mitigate hallucinations in image MLLMs, several strategies have been proposed. Woodpecker (Yin et al. 2023a) introduces a post-hoc correction mechanism leveraging additional visual cues, while Volcano (Lee et al. 2023) adopts a critique-revise-decide framework for iterative self-revision. From a data perspective, LRV-Instruction (Liu et al. 2023a) and HalluciDoctor (Yu et al. 2024) aim to alleviate hallucinations by constructing more balanced, high-quality instruction tuning datasets.
Another line of work addresses hallucinations during inference, where contrastive decoding (CD) has shown effectiveness without requiring additional training. VCD (Leng et al. 2024) mitigates statistical bias and language priors by contrasting original and distorted images during decoding. However, existing CD methods mainly target ImageLLMs, with limited exploration in the video domain. Motivated by this gap, we propose Temporal Contrastive Decoding (TCD) to reduce event hallucinations in VideoLLMs.
@@
Susceptibility to vision-language correlation biases. In this category, the primary challenge lies in sourcing the appropriate video data. Since existing publicly available video datasets (e.g., ActivityNet (Caba Heilbron et al. 2015), (Das et al. 2013), (Soomro 2012)) collect videos from daily life, the events they include cater to the vision-language correlation biases stored in VideoLLMs, resulting in high recognition precision. Therefore, instead of refactoring existing video datasets, we meticulously collect videos from prevalent Internet video platforms and filter them with rigorous cross-validation. Specifically, we begin by manually collecting videos that depict infrequent events based on human consensus. To minimize personal bias, we engage three annotators to assess whether the events in these selected videos are indeed uncommon. Only videos that receive unanimous positive judgments from all annotators are retained as testing samples. Similar to the first category, we also employ the blurring operation on videos to prevent information leakage. We further divide these videos into two groups: (1) Videos where infrequent events occur throughout the entire duration, referred to as “Entire Rare Events” in Fig.2. (2) Videos containing a mix of both frequent and infrequent events, referred to as “Mix Common-Rare Events” in Fig.2.
@@
Binary: “Did/Is the [subject] [event] in the video?”,
@@
“Did the [subject] [event] in the entire video?”,
@@
Evaluation Protocol
Binary question answering. Following previous benchmarks (Wang et al. 2024; Li et al. 2023b), we compare the first word of the model’s answer with the ground-truth answer to determine whether the model answered the question correctly. This approach is justified by the VideoLLM’s exceptional instruction-following capability, which typically enables it to answer binary choice questions with “yes” or “no” at the beginning. We calculate the accuracy as the evaluation metric for binary questions.
Open-ended question answering. For automatic evaluation, we feed both generated and ground-truth answers into GPT-4o(Hurst et al. 2024), using it as a knowledgeable evaluator to determine whether the model answers the question correctly following prior benchmarks (Liu et al. 2024b). We also use accuracy as the evaluation metric. Detailed templates of prompting GPT-4o are provided in the supplementary materials due to page limitation.
@@
VideoChatGPT (Maaz et al. 2023) 14.91 5.50 56.99 3.63 21.57 36.43 4.30
VideoChat2 (Li et al. 2024b) 16.67 4.59 47.67 1.55 22.55 32.76 2.64
PLLaVA (Xu et al. 2024) 45.61 16.51 58.55 3.11 81.37 60.64 6.05
LLaMA-VID (Li, Wang, and Jia 2023) 30.70 16.51 73.58 7.77 43.14 54.03 10.90
LLaVA-NeXT-Video (Li et al. 2024a) 49.12 9.17 70.47 4.66 71.57 64.79 6.29
VILA (Lin et al. 2024) 53.51 20.18 58.55 17.62 83.33 63.33 18.54
ShareGPT4Video (Chen et al. 2024) 11.40 0.00 67.88 5.18 6.86 49.14 9.82
Video-LLaVA (Lin et al. 2023) 30.70 8.26 57.51 7.25 41.18 45.97 7.62
Closed-source Models
GPT-4V (Achiam et al. 2023) 57.89 11.01 79.27 16.58 92.16 76.53 14.57
Gemini-1.5-Pro (Team et al. 2023) 61.40 49.54 83.42 39.90 96.08 80.44 43.38
GPT-4o (Hurst et al. 2024) 83.33 58.72 92.75 41.97 100.00 91.93 48.01
Table 1: Performance comparison of VideoLLMs on EventHallusion. We exhibit the accuracy of both open-source and closed-source models on the binary classification task and the detailed description matching task on our three types of videos, respectively. We highlight the best-performing results for both closed-source and open-source VideoLLMs in bold. In addition, we show the overall success rate of each model on our binary classification task that answers ”yes” or ”no” in the first word.
Temporal Contrastive Decoding
To mitigate the aforementioned hallucinations in VideoLLMs, the most straightforward solution is to enhance finetuning with additional debiased data. However, this approach is highly resource-intensive, as it requires extensive efforts to collect such videos and annotate corresponding QA pairs. Therefore, we take inspiration from (Leng et al. 2024) and propose our Temporal Contrastive Decoding (TCD) method, which alleviates event-oriented hallucinations in VideoLLMs by comparing the logits of the original video and its counterpart during the decoding stage.
@@
where $S(\cdot)$ represents the temporal downsampling operation. With the generated $z_{t}^{ori}$ and $z_{t}^{con}$, the modulated logit at the current timestep $t$ can be calculated as:
@@
where $\alpha$ is the hyperparameter to control the contrastive strength during the decoding. Since $z_{t}^{ori}$ and $z_{t}^{con}$ are not always exhibited contrastively, we eliminate the adverse effects of suppressing true positives by directly filtering implausible answers with a dynamic confidence threshold $t$:
@@
[…]
$p_{t}$ $=\mathrm{softmax}(\tilde{z}_{t}).$
@@
Model Decoding Entire Mix Mis. Overall
LNV (Li et al. 2024a) Default 49.12 70.47 71.57 64.79
@@
Table 2: Results of TCD on EventHallusion. Default denotes the standard decoding strategy, while TCD represents our Temporal Contrastive Decoding method. We report the accuracy of the binary classification task. The best results of each category are bolded. “LNV” and “VC2” are short for “LLaVA-NeXT-Video” and “VideoChat2”, respectively. “Mis” denotes the “Misleading” category.
Experiments
In this section, we first evaluate multiple prevalent VideoLLMs on our EventHallusion, including both open-source and closed-source MLLMs. Subsequently, we examine the effectiveness of our proposed TCD and also discuss its more specific design choices.
Evaluation on EventHallusion
Performances of Human Evaluation
Firstly, to evaluate the quality of our proposed EventHallusion benchmark, we ask three humans to answer the questions and average their resulting scores as the human evaluation results. Meanwhile, to validate the value of evaluating event-oriented hallucinations for VideoLLMs, we also extracted object contents from videos in our EventHallusion benchmark and tested them using both human evaluations and two prevalent VideoLLMs. The above results are shown in Fig.6, where we have the following conclusions: (1) Although prevalent VideoLLMs perform poorly in EventHallusion (i.e., “Event” part in Fig.6), humans can correctly answer almost all questions. This demonstrates that the questions in our EventHallusion benchmark are clear enough to allow for accurate judgments. (2) When recognizing objects in videos, prevalent VideoLLMs can achieve human-level accuracies. However, they significantly lag behind humans when comprehending event-related content. This demonstrates that event-oriented hallucinations confuse VideoLLMs much more severely than object-related ones.
Performances of Prevalent VideoLLMs
Evaluation Setup. We test 8 open-source Video and 3 closed-source VideoLLMs on our EventHallusion. Open-source VideoLLMs comprise VideoChatGPT (Maaz et al. 2023), VideoChat2 (Li et al. 2024b), PLLaVA (Xu et al. 2024), LLaMA-VID (Li, Wang, and Jia 2023), LLaVA-NeXT-Video (Li et al. 2024a), VILA (Lin et al. 2024), ShareGPT4Video (Chen et al. 2024), and Video-LLaVA (Lin et al. 2023). For closed-source VideoLLMs, we select the three most popular video-supported MLLMs, namely GPT-4V (Achiam et al. 2023), GPT-4o(Hurst et al. 2024), and Gemini-1.5-Pro (Team et al. 2023), to conduct the systematic comparison of the robustness between open-source and closed-source MLLMs.
For the inference setting of each VideoLLM, we follow their original setting, including conversation mode and other hyperparameters. For the number of frames, we sample at least 8 frames for each VideoLLM to ensure the complete preservation of event information. In addition, to ensure the stability and reproduction of our results, we adopt the greedy search as the default decoding strategy for all open-source models. For the closed-source models, we sample 8 frames.
@@
Performance Analysis. We exhibit the performance of each VideoLLM in Tab 1 on the binary classification and detailed description matching task. As the results indicate, open-source VideoLLMs exhibit a consistent weakness in event hallucinations. In contrast, closed-source VideoLLMs demonstrate greater robustness against event hallucinations across all categories. A deeper analysis is provided below: (1) In evaluating the susceptibility of VideoLLMs to vision-language correlation biases (i.e., “Entire” and “Mix” in Tab.1), most VideoLLMs perform worse when confronted with videos containing a mix of common and rare events, compared to those containing only rare events. This result indicates that frequent events can distract the model’s attention, imposing higher challenges to the model. (2) In evaluating the susceptibility of VideoLLMs to language priors (i.e., “Misleading” in Tab.1), closed-source models demonstrate overwhelming advantages compared to open-source models, suggesting that these closed-source VideoLLMs are built upon language foundation models that are significantly more powerful than their open-source counterparts.
TCD Evaluation
VideoLLM Baselines.
To evaluate the effectiveness of our TCD in addressing event hallucinations, we adopt LLaVA-NeXT-Video (Li et al. 2024a), VideoChat2 (Li et al. 2024b), and VILA (Lin et al. 2024) as the foundation backbones. LLaVA-NeXT-Video extends the ability of LLaVA (Liu et al. 2024a) based on the projection layer design, while VideoChat2 adopts the Q-Former (Li et al. 2023a) to bridge the gap between videos and LLM. VILA, on the other hand, employs enhanced visual-language pre-training for better video comprehension. Based on these MLLMs, we explore the performance of TCD across different VideoLLM architectures and designs on EventHallusion.